# Effect of freestream turbulence on the coherent dynamics of a wind turbine wake

### v1

@@ 5 Modal decomposition
The coherent motions present in a wind turbine wake have been studied using different modal decomposition techniques such as Proper Orthogonal Decomposition (POD), Dynamic Mode Decomposition (DMD) or some variant of the two [Sarmast et al., 2014, Debnath et al., 2017, De Cillis et al., 2021, Biswas and Buxton, 2024b]. These methods successfully captured the modes related to tip vortices and vortex sheddings from the tower or nacelle and wake meandering [Debnath et al., 2017, De Cillis et al., 2021, Kinjangi and Foti, 2023, Biswas and Buxton, 2024b]. For the current dataset, we use Optimal Mode Decompoisition (OMD) which is a more generalised version of the originally proposed DMD algorithm [Schmid, 2010, Wynn et al., 2013]. A detailed description of the OMD based triple decomposition technique can be found in our previous works [Baj et al., 2015, Baj and Buxton, 2017, Biswas et al., 2022, Biswas and Buxton, 2024b]. Biswas and Buxton [2024b] recently applied OMD on the wake produced from the same wind turbine model as in the present study. Various coherent modes such as the ones associated with the tip vortices, the vortex sheddings from the nacelle and the tower and wake meandering were identified and the effect of the tip speed ratio was discussed. However, the analysis was restricted to the case without any freestream turbulence. In this section, we build on the results of Biswas and Buxton [2024b] by including the effect of freestream turbulence.
@@ 5.1 Mode clustering
As discussed previously, the nature of the power spectra in the vicinity of the wake meandering frequency band (WMFB) i.e. in $0.15\leq St_{D}\leq 0.4$, changed from a more discrete nature to a more broadband nature when $T_{i}$ was increased. Due to this broadband nature of the wake meandering in the presence of FST, when OMD was performed, the FST cases showed a comparatively larger number of modes in the WMFB for which the mode shapes looked physically meaningful. An example OMD spectrum for case 2a is shown in figure 10(a). It was therefore imperative to group the spatially similar modes together and to represent them by a single representative mode. Beit-Sadi et al. [2021] recently proposed a clustering algorithm that grouped/clustered similar OMD modes together to form a smaller set of dissimilar modes. The authors used graph theory to cluster the modes in an iterative process. To assess the similarity of two modes, the authors looked at the similarity between the subspaces spanned by the real and imaginary parts of the complex OMD modes. The algorithm was later successfully applied to construct the shedding modes behind a cylinder exposed to FST [de Oliveira et al., 2025]. The procedure adopted here differs slightly from the original algorithm. The original algorithm is briefly discussed here along with the changes introduced. For each OMD mode $\phi_{l}$, Beit-Sadi et al. [2021] defined the metric $A_{l}=[\Re(\phi_{l}),\Im(\phi_{l})]$, where $\Re$ and $\Im$ denote the real and imaginary parts of $\phi_{l}$ and are stored as column vectors. To assess the similarity between modes the authors used the smallest angle between the subspaces of the modes. The smallest angle between two modes $\phi_{l}$ and $\phi_{m}$ was therefore measured by using the statistic $\theta_{lm}=\sin^{-1}\Big{(}{||A_{l}-A_{m}\Big{(}A_{l}^{T}A_{m}\Big{)}||}_{2}\Big{)}$, where ||.||||.||| | . | | denoted the largest singular value of the matrix. For simplicity, the spatial similarity was then represented by $\zeta=\cos(\theta_{lm})$.
Figure 10: $1^{st}$, $2^{nd}$ and $5^{th}$ iterations of the mode clustering algorithm for the case 2a. The chosen parameters were as follows: $r=175$, $\zeta_{spat}=0.94$, and $\zeta_{spec}=0.3$.
A spatial similarity cutoff ($\zeta_{spat}$) was defined, above which two modes could be considered similar, or dissimilar otherwise. As an example, for the spectrum shown in figure 10, clustering is performed with $\zeta_{spat}=0.94$ (note that $\zeta_{spat}=1$ means the two modes span the same subspace). This $\zeta_{spat}$ is selected based on a convergence analysis which will be discussed in the subsequent section. Additionally, a spectral similarity cutoff $\zeta_{spec}$ was imposed such that two spatially similar modes would not be clustered if their Strouhal numbers were $\zeta_{spec}$ (or more) apart. In figure 10, $\zeta_{spec}$ was set to 0.3 which is of the order of the bandwidth of the WMFB. Nonetheless, the clustering process only weakly depended on $\zeta_{spec}$ as will be shown. The spatially similar modes are then connected with an ‘edge’ (shown by the blue lines). Note from figure 10(a) that there exists a number of similar modes only in the vicinity of the WMFB. Once the similar modes are identified, the problem of clustering becomes equivalent to finding subgroups where each mode is connected with every other mode with an edge. In graph theoretic notion, such a subgroup is termed as a ‘clique’. Further, a ‘maximal-clique’ is a subgroup or clique that is not strictly contained in any other cliques (see Beit-Sadi et al. [2021] for more details).
Once the maximal cliques are obtained, each maximal clique is represented using a single representative mode. For that, each of the modes in a maximal clique are stacked in a matrix (let us denote it as $M$) as column vectors multiplied by a significance parameter $\sigma_{l}$. If the $k^{th}$ maximal clique has $n$ modes, then $M_{k}=[\sigma_{1}\phi_{1},\sigma_{2}\phi_{2}...\sigma_{n}\phi_{n}]$. Thereafter, a singular value decomposition of $M_{k}$ is performed and the first singular vector is used as the mode representing the cluster. However, it is worthwhile to note that several aspects of the method are still not strictly defined and deserve further consideration. One of them is how to assign a frequency and growth rate to a clustered mode. Further, which statistic is the best to use as $\sigma$? For example, one could project the snapshots on the OMD modes to obtain the time varying coefficients of the modes. The magnitude of the coefficient at a specific time or a norm of the coefficient could be used as $\sigma$. Here however, it was found that using such a statistic undermined the influence of the lesser damped modes within a cluster. Accordingly, a quantity related to the growth rate ($\Re(\lambda_{OMD})$) of the modes was thought to be a more reasonable measure of $\sigma$ motivating the use of the quantity $1/|\Re(\lambda_{OMD})|$. The dominant modes (growth rate closer to zero) are therefore assigned a higher significance value. This method showed faster convergence for the current data set and yielded better converged modes representing a cluster. The modes representing a cluster are also assigned a virtual growth rate and a frequency which is obtained by taking a weighted average of the frequencies and growth rates of the modes forming a cluster with $\sigma$ used as the weight. The whole process is repeated until no two modes are similar. The $1^{st}$, $2^{nd}$ and the $5^{th}$ (final) iterations for the case 2a are shown in figure 10(a-c). Note that the mode clustering algorithm significantly simplifies the spectrum in the vicinity of WMFB.
Figure 11: Variation of the number of converged modes ($N_{c}$) with $\zeta_{spat}$ for different $\zeta_{spec}$ for the cases (a) 2a, (b) 3 and (c) 4b.
The choice of the cutoff parameters $\zeta_{spat}$ and $\zeta_{spec}$ also involves some subjectivity. The effects of $\zeta_{spat}$ and $\zeta_{spec}$ on the converged number of modes ($N_{c}$) for three different cases are shown in figure 11. The rank of OMD was set to 175 (i.e. the total number of initial modes $N=175$) for all cases following our previous work [Biswas and Buxton, 2024b]. No mode reduction is achieved for $\zeta_{spat}\geq 0.96$, while for $\zeta_{spat}\leq 0.9$ the spectra are over simplified leaving too few modes. A different $\zeta_{spat}$ was selected for each case such that $N_{c}/N\approx 0.9$. For the three cases shown in figure 11, the selected cutoffs are shown by the vertical dashed line. The different symbols in figure 11 correspond to different values of $\zeta_{spec}$. Note that the mode clustering process does not depend significantly on $\zeta_{spec}$ for $\zeta_{spec}\geq 0.2$. The analysis was therefore continued with $\zeta_{spec}=0.3$ for all the cases.
Figure 12: $1^{st}$, $2^{nd}$ and final iterations of the mode clustering algorithm for the no grid case (a-c), case 3 (d-f) and case 4b (g-i).
@@
The first, second and the final iterations for the no grid case, case 3 and case 4b are shown in figures 12. Interestingly, a significant number of similar modes are observed for the no grid case as well. However, the cluster contained a mode (with $St_{D}\approx 0.22$) that was significantly less damped compared to the other modes. This is more clearly observed in figure 12(b). Note that, the recent work by Biswas and Buxton [2024b] applied OMD on the same data set as case 1 here without mode clustering. The modes obtained with/without mode clustering were similar for the no grid case as will be discussed in the next paragraph. Mode clustering therefore simplified the spectrum in the present case, making it easier to select the physically meaningful modes which are used to form a reduced order representation of the coherent component of the flow. Further details about the mode selection process can be found in appendix 2.
Upon clustering, the selected modes are highlighted in figure 12(c) with the ‘o’ symbols. Among the tip vortex related modes, only the frequencies $f_{r}-4f_{r}$ could be captured in the OMD spectrum (figure 12(c)). The modes corresponding to $5f_{r}$ and $6f_{r}$ were therefore obtained using phase averaging, similarly to the method discussed in our previous works [Baj and Buxton, 2017, Biswas and Buxton, 2024b]. At first glance, the number of similar modes in the first iteration for the cases with FST look comparable to that for the no grid case. However, in general, equal or more iterations were required for the cases with FST to reach convergence. Furthermore, for the cases with an elevated level of FST, the clusters were not dominated by a single mode unlike case 1. For all cases, the final modes (marked by the ‘o’ symbols) were carefully selected by looking at the spatial mode shape and how the clusters evolved over iterations.
@@ 5.2 The nature of the wake meandering mode in the presence of FST
Two dominant wake meandering modes (having a frequency in the WMFB) for the four cases of increasing $T_{i}$ are shown in figure 13. For the no grid case (figure 13(a,b)), regular patterned structures of wavelength $1.5-2D$ are observed. Note that, these two modes are almost exactly the same as the modes reported by Biswas and Buxton [2024b] without mode clustering (see figures 4(a,b)), apart from a slight phase shift and a slight difference in the mode amplitude. This offers reassurance that the mode clustering algorithm adopted here does not alter the physical information embedded in the data set, rather it simplifies the analysis. For case 2a (low $T_{i}$), the dominant wake meandering modes in figures 13(c,d) still show similar regular structures, but the modes exhibit a larger transverse (along $y$) width compared to that for the no grid case. Note from figure 13(a,b) that the modes for the no grid case are not particularly energetic for $y\geq 0.5D$. The modes initiate from the vicinity of the nacelle highlighting the importance of the nacelle to providing the perturbation necessary to kick start wake meandering [Howard et al., 2015, Biswas and Buxton, 2024a]. For higher $T_{i}$, the modes show a significant energy content for $y\geq 0.5D$ as was discussed in section 4.2. The early breakdown of the tip vortices in the presence of elevated FST can provide additional perturbations in the vicinity of the outer wake. As $T_{i}$ is increased further, the modes become more irregular in nature. A higher transverse width of the modes can however still be observed, at least for the medium $T_{i}$ case (case 3).
@@
Figure 13: Two dominant wake meandering modes for the cases (a,b) without FST, (c,d) case 2a, (e,f) case 3, (g,h) case 4b.
Figure 14: Modes associated with $3f_{r}$ (a,c,e,g) and $f_{r}$ (b,d,f,h) for the no grid case (a,b), case 2a (c,d), case 3 (e,f) and case 4b (g,h).
@@ 5.3 The modes related to the tip vortices
Among the tip vortex related frequencies, only the modes associated with $3f_{r}$ and $f_{r}$ are shown in figure 14 as they were the dominant near wake structures. The modes associated with $3f_{r}$ clearly show how the tip vortices break down earlier as freestream turbulence is increased. A similar pattern is observed for the modes associated with $f_{r}$ as well as the harmonics of $f_{r}$ and $3f_{r}$ (not shown for brevity). Furthermore, the modes with FST appear to have a larger transverse spread. This is likely due to the vortex wandering or vortex jittering reported in previous studies in the presence of FST [Gambuzza and Ganapathisubramani, 2023]. Note that for cases 2a and 2b, new frequencies were observed that corresponded to the sum and difference of $3f_{r}$ and a wake meandering frequency that was dominant near the edge of the wake. These frequencies were however not present in the OMD spectrum for either of the cases. A similar observation was reported in our previous studies where the new frequencies corresponding to the sum and difference of the dominant coherent modes were not captured by OMD [Baj and Buxton, 2017, Biswas et al., 2022]. These modes were therefore obtained using phase averaging as discussed in Baj and Buxton [2017]. These are shown in figure 15 for cases 2a and 2b. The spatial orientation of the modes reveals an interesting pattern. Note from figure 6(b) that the $f_{wm}$ mode interacting with $3f_{r}$ was energetic in the freestream side ($y>0.5D$). In order to indicate the difference in the spatial location of the modes, a black line was added to figure 15 showing the loci of the points where the kinetic energy of the mode $3f_{r}$ was maximum. It can be seen that for both the cases, the high frequency mode $3f_{r}+f_{wm}$ is situated more towards the freestream side, while $3f_{r}-f_{wm}$ resided more towards the wake side. The exact same spatial orientation of the frequencies $f_{c}\pm f_{m}$ was reported by Biswas et al. [2022], Baj and Buxton [2017] for fundamentally different multiscale flows. Here, the high frequency mode resided more towards the shear layer between the two wake generating bodies, while the low frequency mode resided in the other outer side of the wake of the smaller wake generating body (having a higher shedding frequency). Such similarity between a simple two dimensional multiscale flow and a complex three dimensional multiscale flow involving rotation is truly remarkable and adds credance to the notion that these multiscale interactions are indeed universal.
Figure 16: Integrated budget terms for (a) no grid case, (b) case 2a, (c) case 3 and (d) case 4b.
@@ 6 Energy exchanges
In this section, we use the multi-scale triple-decomposed coherent kinetic energy budget equation derived by Baj and Buxton [2017] to assess the energy budgets of the individual coherent modes. Biswas and Buxton [2024b] recently used this equation to obtain the energy budget terms of each of the individual coherent modes in a wind turbine wake in the absence of FST. It was reported that the modes associated with the nacelle/tower’s vortex shedding and wake meandering were energised by the mean flow. Contrastingly, the tip vortex related modes could be energised through different pathways such as energy transfer from the mean flow or non-linear triadic interactions between various coherent modes co-existing in the wake. The role of the traidic interactions in the tip vortex stability and merging process was quantified for the first time. Yeung et al. [2024] recently proposed a Triadic Orthogonal Decomposition (TOD) which identified coherent structures that optimally account for triadic interactions and quantified the triadic energy transfers. The authors used the same PIV dataset obtained by Biswas and Buxton [2024b] and reported that the energy transfers obtained using the new method were in good agreement with Biswas and Buxton [2024b]. Although these studies have significantly improved our understanding of the origin and evolution of the various coherent structures in a wind turbine wake, the effect of freestream turbulence was not considered in these studies. This section attempts to fill that gap.
The triple-decomposed coherent kinetic energy budget equation derived by Baj and Buxton [2017] can be symbolically represented as the following:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\partial\tilde{k}_{l}}{\partial t}=-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{C}_{l}}+\tilde{P}_{l}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{P_{l}}}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Big{(}\tilde{T}^{+}_{l}-\tilde{T}^{-}_{l}\Big{)}}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{\epsilon_{l}}}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{D_{l}}}}$ (1)
Here, the source terms on the right hand side consist of convection ($\tilde{C}_{l}$), mean flow production ($\tilde{P}_{l}$) or energy transfer from the mean flow to the $l_{th}$ coherent mode, production of stochastic turbulent kinetic energy directly from the coherent mode $l$ ($\hat{P}_{l}$), triadic energy production ($\tilde{T}^{+}_{l}-\tilde{T}^{-}_{l}$), direct dissipation from coherent mode $l$ (${\tilde{\epsilon}}_{l}$) and diffusion ($\tilde{D}_{l}$). Baj and Buxton [2017] showed that the triadic energy production term can become significant only when there exists three frequencies that linearly combine to zero or, in other words, there is a triad in the form $f_{l}\pm f_{m}\pm f_{n}=0$. The constituents of this term were defined as follows:
@@
Here, $\tilde{T}^{+}_{l}$ denotes the nonlinear energy transfer from $f_{s}$ to $f_{l}$, while $\tilde{T}^{-}_{l}$ denotes the energy lost by $f_{l}$ to $f_{s}$. In both these terms, $f_{t}$ acts as a mediator or a catalyst [Baj and Buxton, 2017, Yeung et al., 2024]. The total triadic energy gain/loss for $f_{l}$ is obtained by summing over all the frequencies ($f_{s}$, $f_{t}$) that form a triad with $f_{l}$. Note that multiple such triads can be readily identified in the tip vortex system of a wind turbine wake. Biswas and Buxton [2024b] reported these triadic interactions in the tip vortex system for a range of $\lambda$s and quantified the importance of this term on the tip vortex merging process. The other term they found to be of particular interest was the $\tilde{P}_{l}$ term since it was directly linked to wake recovery. $\tilde{P}_{l}$ is defined as $\tilde{P}_{l}=-\sum_{m}\overline{\tilde{u}^{m}_{i}\tilde{u}^{l}_{j}}\frac{\partial\overline{u}_{i}}{\partial x_{j}}$, where $m$ can be any coherent mode selected in the reduced order representation of the flow including the $l_{th}$ mode. The complete composition of each of the terms of equation 1 can be found in Baj and Buxton [2017].
@@
Figure 17: Profiles of normalised coherent velocity correlation ($\overline{\tilde{u}^{l}\tilde{v}^{l}}$) and transverse gradient of mean streamwise velocity ($\frac{\partial\overline{u}}{\partial y}$) for the wake meandering mode extracting the highest amount of energy from the mean flow for the no grid case (a, b) and case 2a (c, d). The profiles are shown at $x/D=1$ (a, c) and $x/D=3$ (b, d).

### v2

@@ 4 Modal decomposition
The coherent motions present in a wind turbine wake have been studied using different modal decomposition techniques such as Proper Orthogonal Decomposition (POD), Dynamic Mode Decomposition (DMD) or some variant of the two [Sarmast et al., 2014, Debnath et al., 2017, De Cillis et al., 2021, Biswas and Buxton, 2024b]. These methods successfully captured the modes related to tip vortices and vortex sheddings from the tower or nacelle and wake meandering [Debnath et al., 2017, De Cillis et al., 2021, Kinjangi and Foti, 2023, Biswas and Buxton, 2024b]. For the current dataset, we use Optimal Mode Decompoisition (OMD) which is a more generalised version of the originally proposed DMD algorithm [Schmid, 2010, Wynn et al., 2013]. DMD assumes that the time evolution of the flow can be governed by a time invariant, best-fit linear operator $\boldsymbol{A}$ and the eigen-decomposition of the operator gives the so-called DMD modes. Here the time dynamics are obtained by projecting the flow data onto a POD subspace. Contrastingly, in OMD, a more generalized approach is followed by solving a two-way optimization problem for the a low order subspace of the flow and it’s dynamics. Wynn et al. [2013] showed that the restrictive nature of DMD can result into significant under performance in some cases compared to OMD [Wynn et al., 2013]. In our previous work, we compared the performance of OMD with POD in a different flow considering the wake of a multiscale array of bars [Baj et al., 2015]. POD was found to be incapable of correctly identifying the mode shapes further downstream from the wake generating body. Therefore, OMD was the preferred choice. A brief overview of OMD is presented in appendix 2.
Biswas and Buxton [2024b] recently applied OMD to the wake produced by the same wind turbine model as in the present study. Various coherent modes such as the ones associated with the tip vortices, the vortex sheddings from the nacelle and the tower and wake meandering were identified and the effect of the tip speed ratio was discussed. However, the analysis was restricted to the case with negligible freestream turbulence. In this section, we build on the results of Biswas and Buxton [2024b] by including the effect of freestream turbulence.
@@ 4.1 Mode clustering
OMD was performed with a rank $r=175$ to be consistent with our previous work [Biswas and Buxton, 2024b]. For brevity, the typical OMD spectra are only shown for four selected cases: 1 (no grid case), 2c, 3b and 4b in figures 16(a, d, g, j). A mode with a higher growth rate ($\Re(\lambda_{OMD})$) generally represents a physically meaningful mode. Due to the discrete nature of the tip vortex related frequencies, their identification from the OMD spectra was straightforward. However, it was not so for the low frequency wake meandering related modes. In the vicinity of the wake meandering frequency band (WMFB) i.e. in $0.15\leq St_{D}\leq 0.4$, there were multiple modes with comparable frequencies and growth rates. The associated mode shapes also looked similar. This was more of a problem for the cases with medium to high FST levels. Therefore, in order to simplify the OMD spectra (especially in the vicinity of the WMFB), the spatially (and spectrally) similar modes were identified and grouped together into single representative mode. Beit-Sadi et al. [2021] proposed a clustering algorithm that grouped/clustered similar OMD modes together to form a smaller set of dissimilar modes. The authors used graph theory to cluster the modes in an iterative process. To assess the similarity of two modes, the authors looked at the similarity between the subspaces spanned by the real and imaginary parts of the complex OMD modes. A constraint based on spectral similarity was also imposed. Finally, two modes were considered to be similar if the spatial and spectral similarity was above a threshold value. The algorithm was later successfully applied to construct the shedding modes behind a cylinder exposed to FST [de Oliveira et al., 2025]. We apply the same algorithm for this case, albeit with some changes in the way the similar modes are grouped. The method is discussed in detail in appendix 2.
The evolution of the OMD spectra at the first, second and the final iterations of clustering is shown in figure 16 for the four selected cases. For each case, a sensitivity analysis was performed to select a threshold value/cutoff for spatial and spectral similarity (denoted by $\zeta_{spat}$ and $\zeta_{spec}$ respectively). This is discussed in appendix 2 (please see figure 25). After finding similar modes, they are connected with an ‘edge’ (shown by the black lines). Note that there exists a number of similar modes only in the vicinity of the WMFB. Once the similar modes are identified, the problem of clustering becomes equivalent to finding subgroups/clusters where each mode is connected with every other mode with an edge. In addition, a subgroup should not be strictly contained in another group. In order to obtain the mode representing a cluster, the modes in the same cluster are multiplied with a weight (related to the growth rate of the modes) and stored in a matrix. Thereafter, the first singular value of the matrix was simply used to represent the mode associated to the cluster. Please refer to appendix 2 for further details.
@@
Interestingly, a significant number of similar modes are observed for the no grid case as well. Note that, the recent work by Biswas and Buxton [2024b] applied OMD on the same data set as case 1 here without mode clustering. The modes obtained with/without mode clustering were similar for the no grid case as will be discussed in the next paragraph. Mode clustering therefore simplified the spectrum in the present case, making it easier to select the physically meaningful modes which are used to form a reduced order representation of the coherent component of the flow. Further details about the mode selection process can be found in appendix 3.
Figure 17: Two dominant wake meandering modes for the cases (a,b) 1, (c,d) 2c, (e,f) 3b, (g,h) 4b. The associated Strouhal numbers ($St_{U_{b},D}$) are also highlighted in the figures.
[…]
@@ 4.2 The nature of the wake meandering mode in the presence of FST
The transverse velocity component ($\tilde{v}_{l}$, where $l$ runs over the coherent modes) of the two dominant wake meandering modes (having a frequency in the WMFB) for the four cases of increasing $T_{i}$ are shown in figure 17. For the no grid case (figure 17(a,b)), regular patterned structures of wavelength 1.52D1.5-2D are observed. The modes start from close to the nacelle in the inner part of the wake and expand further downstream. Note that, these two modes are almost exactly the same as the modes reported by Biswas and Buxton [2024b] without mode clustering (see figures 4(a,b)), apart from a slight phase shift and a slight difference in the mode amplitude. This offers reassurance that the mode clustering algorithm adopted here does not alter the physical information embedded in the data set, rather it simplifies the analysis. For case 2c (low $T_{i}$), the dominant wake meandering modes in figures 17(c,d) still show similar regular structures, but the modes exhibit a larger transverse (along $y$) width compared to that for the no grid case. Note from figure 17(a,b) that the modes for the no grid case are not particularly energetic for $y\geq 0.5D$. The modes initiate from the vicinity of the nacelle highlighting the importance of the nacelle to providing the perturbation necessary to kick start wake meandering [Howard et al., 2015, Biswas and Buxton, 2024a]. For higher $T_{i}$, the modes show a significant energy content for $y\geq 0.5D$ as was discussed in section 3. They also become increasingly irregular in nature with increasing FST levels. A higher transverse width of the modes can however still be observed, at least for the medium $T_{i}$ case (case 3b).
@@
Figure 18: Modes associated with $3f_{r}$ (a,c,e,g) and $f_{r}$ (b,d,f,h) for the no grid case (a,b), case 2c (c,d), case 3b (e,f) and case 4b (g,h).
@@ 4.3 The modes related to the tip vortices
Among the tip vortex related frequencies, only the modes associated with $3f_{r}$ and $f_{r}$ are shown in figure 18 for the same four cases (1, 2c, 3b, 4b) since they were the most important and were observed for all the cases (see figure 7). The modes associated with $3f_{r}$ clearly show how the tip vortices break down earlier as freestream turbulence intensity is increased. A similar pattern is observed for the modes associated with $f_{r}$ as well as the harmonics of $f_{r}$ and $3f_{r}$ (not shown for brevity). Furthermore, the modes with FST appear to have a larger transverse spread. This is likely due to the vortex wandering or vortex jittering reported in previous studies in the presence of FST [Gambuzza and Ganapathisubramani, 2023].
@@
Note that for the cases 2a and 2b, new frequencies were observed that corresponded to the sum and difference of $3f_{r}$ and a wake meandering frequency that was dominant near the edge of the wake. These frequencies were however not present in the OMD spectrum for either of the cases. A similar observation was reported in our previous studies where the new frequencies corresponding to the sum and difference of the dominant coherent modes were not captured by OMD [Baj and Buxton, 2017, Biswas et al., 2022]. These modes therefore had to be obtained using phase averaging. These are shown in figure 19 for cases 2a and 2b. The spatial orientation of the modes reveals an interesting pattern. Note from figure 11(b) that the $f_{wm}$ mode interacting with $3f_{r}$ was energetic in the freestream side ($y>0.5D$). In order to indicate the difference in the spatial location of the modes, a black line was added to figure 19 showing the loci of the points where the kinetic energy of the mode $3f_{r}$ was maximum. It can be seen that for both the cases, the high frequency mode $3f_{r}+f_{wm}$ is situated more towards the freestream side, while $3f_{r}-f_{wm}$ resided more towards the wake side. The exact same spatial orientation was observed for secondary frequencies in fundamentally different multiscale flows [Biswas et al., 2022, Baj and Buxton, 2017]. Biswas et al. [2022] considered the wake of a cylinder (main cylinder) and control rod (another cylinder of diameter 1/10th1/10{th} of the main cylinder). Apart from the shedding modes associated with the main cylinder ($f_{m}$) and control rod ($f_{c}$), secondary frequencies $f_{c}\pm f_{m}$ were observed. The high frequency mode resided more towards the shear layer between the two wake generating bodies, while the low frequency mode resided in the other outer side of the wake of the smaller wake generating body (having a higher shedding frequency). A similar observation was reported by Baj and Buxton [2017] in the wake of a multiscale array of prisms. These secondary frequencies were shown to be a result of non-linear traidic interaction between the two primary shedding frequencies. The fact that we observe such similarity between simple two dimensional multiscale flows and a complex three dimensional multiscale flow involving rotation is truly remarkable and adds credance to the notion that these multiscale interactions are indeed universal.
Figure 20: Summed budget terms for (a) no grid case, (b) case 2c, (c) case 3b and (d) case 4b. The numbers on the right show the associated Strouhal numbers ($St_{U_{b},D}$) of the modes. For the clustered modes, the $St_{U_{b},D}$ values are highlighted in purple. $E$ is a generic variable representing the budget terms.
@@ 5 Energy exchanges
In this section, we use the multi-scale triple-decomposed coherent kinetic energy budget equation derived by Baj and Buxton [2017] to assess the energy budgets of the individual coherent modes. Baj and Buxton [2017] derived separate equations for the kinetic energy associated with the mean flow ($\bar{k}$), each of the coherent modes ($\tilde{k}_{l}$ for the $l_{th}$ coherent mode) and the stochastic component of the flow ($k^{{}^{\prime\prime}}$). For the present work, we only use the equation for the coherent modes which can be symbolically represented as the following:
[…]
Here, the source terms on the right hand side consist of convection ($\tilde{C}_{l}$), mean flow production ($\tilde{P}_{l}$) or energy transfer from the mean flow to the $l_{th}$ coherent mode, production of stochastic turbulent kinetic energy directly from the coherent mode $l$ ($\hat{P}_{l}$), triadic energy production ($\tilde{T}^{+}_{l}-\tilde{T}^{-}_{l}$), direct dissipation from coherent mode $l$ (${\tilde{\epsilon}}_{l}$) and diffusion ($\tilde{D}_{l}$). Note that all the terms in equation 2 are averaged over time. Baj and Buxton [2017] showed that the triadic energy production term can become significant only when there exist three frequencies that linearly combine to zero or, in other words, there is a triad in the form $f_{l}\pm f_{s}\pm f_{t}=0$. The constituents of this term were defined as follows:
@@
Here, $\tilde{T}^{+}_{l}$ denotes the nonlinear energy transfer from $f_{s}$ to $f_{l}$, while $\tilde{T}^{-}_{l}$ denotes the energy lost by $f_{l}$ to $f_{s}$. In both these terms, $f_{t}$ acts as a mediator or a catalyst [Baj and Buxton, 2017, Yeung et al., 2024]. $\tilde{u}^{f_{l}}_{i}$ denotes the $i_{th}$ velocity component of the coherent mode $l$ (having frequency $f_{l}$). The total triadic energy gain/loss for $f_{l}$ is obtained by summing over all the frequencies ($f_{s}$, $f_{t}$) that form a triad with $f_{l}$. The other term that is of particular interest is the $\tilde{P}_{l}$ term since it is directly linked to wake recovery. $\tilde{P}_{l}$ is defined as $\tilde{P}_{l}=-\sum_{m}\overline{\tilde{u}^{m}_{i}\tilde{u}^{l}_{j}}\frac{\partial\overline{u}_{i}}{\partial x_{j}}$, where $m$ can be any coherent mode selected in the reduced order representation of the flow including the $l_{th}$ mode. The complete composition of each of the terms of equation 2 can be found in Baj and Buxton [2017].
Biswas and Buxton [2024b] recently used this equation to obtain the energy budget terms of each of the individual coherent modes in a wind turbine wake with negligible freestream turbulence. It was reported that the modes associated with the nacelle/tower’s vortex shedding and wake meandering were energised by the mean flow. Contrastingly, the tip vortex related modes could be energised through different pathways such as energy transfer from the mean flow or non-linear triadic interactions between various coherent modes co-existing in the wake. The role of the triadic interactions in the tip vortex stability and merging process was quantified for the first time. Yeung et al. [2024] recently proposed a Triadic Orthogonal Decomposition (TOD) which identified coherent structures that optimally account for triadic interactions and quantified the triadic energy transfers. The authors used the same PIV dataset obtained by Biswas and Buxton [2024b] and reported that the energy transfers obtained using the new method were in good agreement with the previous work. Although these studies have significantly improved our understanding of the origin and evolution of the various coherent structures in a wind turbine wake, the effect of freestream turbulence was not considered in these studies. This section attempts to fill that gap.
@@
Figure 21: Profiles of normalised coherent velocity ($\overline{\tilde{u}^{l}\tilde{v}^{l}}$) and transverse gradient of mean streamwise velocity ($\frac{\partial\overline{u}}{\partial y}$) for the wake meandering modes extracting the highest amount of energy from the mean flow for the no grid case (a, e), case 2c (b, f), case 3b (c, g) and case 4b (d,h). The profiles are shown at $x/D=1$ (a-d) and $x/D=4$ (e-h).
The kinetic energy budget terms are first discussed for the four cases analysed previously. The energy budget terms are summed over the entire domain and shown in figure 20. Since the modes related to the tip vortices show weaker energy exchanges compared to the low frequency modes, a separate colour bar is used for them. On the right, the numbers show the Strouhal numbers ($St_{U_{b},D}$) associated with the modes. The modes are also classified as either related to wake meandering ($0.15\leq f_{wm}\leq 0.4$), nacelle’s shedding ($0.4\leq f_{n}\leq 0.6$) or the tower’s shedding ($0.6\leq f_{T}\leq 1$) as discussed in section 3.3. On some occasion this classification could not be made strictly based on the frequency, particularly for the clustered modes. As already discussed, for these modes, the frequency associated with the cluster was obtained as a weighted average of the frequencies of the modes forming the cluster (which may differ from the actual frequency of the mode). Therefore, the clustered modes were carefully classified by looking at their mode shape and the frequency content in the associated time varying mode coefficient. An example can be seen in figure 20(c), where $St_{U_{b},D}=0.1$ was classified as related to wake meandering.